Journalist Report – April 4th

Report title: Journalist Report
Crew #: 333
Position: Crew Journalist
Report prepared by: Joanna Galloway
Date: 04-04-2026
Sol: 6

Journalist Report Title (if applicable): Sol 6

Marie, Joanna and Antoine went out at 9.30am to scout the terrain at Somerville Overview. Once they laid eyes on the rolling landscape, they couldn’t contain their excitement. The terrain never felt so Martian, they could see the desert go on for miles and miles. It was impressive. After a short journey back to the station, they were greeted by rice and chicken salad, not an unknown dish to this crew during this journey. Zahraa and Matthias had been working on their experiments while Matias made lunch. The devastating news finally came out, it seemed like our greenhab had been devastated by some sort of martian rodent.

After lunch, it was time for Joanna and Matias to teach Antoine some boxing defensive moves before the boys had to head back out into EVA. They completed the first EVA by foot since our arrival by going to North Ridge and kept going to Kissing Camel Ridge West. It was a beautiful sight, like most of the views around here. They got to see the station from an above view which showed just how small we are compared to the vast martian views. They decided to get some cardio in and run the way back to the station. A feat considering the heavy equipment on them. They came back feeling pretty proud of themselves. Their EVA finished earlier than expected which allowed them to take a breath before preparing dinner.

Joanna and Marie alternated between cycling and pilates while Zahraa rested. Matthias was still working on his experiment in the greenhab and trying to recuperate what he could from the ravaged legumes. Since tomorrow is our day off, we are planning a movie night with burgers and fries which we are all very excited about!

Sol Summary – April 4th

Report title: Sol Summary Report
Crew #: 333
Position: Crew Commander
Report prepared by: Marie Jansen
Date: 04-04-2026
Sol: 6

Summary Title: Discovery of two new locations for the experiments.

Mission Status: Going great, half way through

Sol Activity Summary:
We woke up at 8:00 AM and had pancakes for breakfast. Antoine, Joanna and Marie then got ready for an EVA. They headed towards Somerville Overview. Once at Compass Rock they had some difficulty finding the route but eventually managed. They loved the view from there but unfortunately the location is too far for Joanna’s experiment. They then headed back with a little stress as the battery of Rover Opportunity dropped very quickly all of a sudden. However they managed to return with both rovers. The crew had a rice salad with chicken for lunch along with tomatoes from the GreenHab. Matias also discovered that an animal (Martian) is eating the GreenHab plants. In the afternoon, Antoine and Matias went on an EVA towards North Ridge and Kissing Camel Ridge West. They headed out on foot to test the range of the walkie-talkies. Despite some communication issues, everything went well and they really enjoyed the view. For the evening, we have planned a little burger night with games, and tomorrow is a rest day for Easter.

Look Ahead Plan:
Tomorrow we are celebrating Easter so we have decided not to work for the day.

Anomalies in work:
The space groundhog eating our vegetables.

Weather:
The weather was really nice today.

Crew Physical Status:
Everyone is doing well.

EVA:
Both EVAs went well! We had a great day today.

Reports to be filed:

  • Sol Summary
  • Operations Report
  • Greenhab Report
  • Journalist Report
  • Astronomy Report
  • EVA Report
  • Daily photos

Support Requested: None

EVA Report – April 4th

Report title: EVA Report
Crew #: 333
Position: Astronomer
Report prepared by: Matias Ballivian Ascarrunz
Date: 04-04-2026
Sol: 6

EVA #: 9

Purpose of EVA: Joanna will conduct this EVA to find the terrain to carry out her experiment next week. Antoine will also make the most of this time and see if the terrain corresponds to his need for a further experiment next week. In the meantime, Marie will give the needed feedback on the terrain for both experiments.

Start time: 8:38 AM
End time: 12:22 PM

Narrative: They started decompression at 8:33 AM, went to their rover and started the way to Somerville Overview. Crew arrived at 9:56 AM to Compass Rock where they didn’t find the correct route directly. They corrected the course and arrived at Somerville Overview at 10:10 AM. Where they explore the zone and take photos for the analysis of the terrain for their experiences. After exploring for a bit they went back to their rovers and got back to the station at 12:15 PM and decompressed before getting back in the HAB.

Destination per MDRS Map: Somerville Overview (and if enough time Watney road)
Destination Coordinates (use UTM WGS 84): 522000E, 4253000N (and 520000E, 4250500N)

EVA Participants:
Joanna, Journalist (Team lead)
Marie, Commander
Antoine, Executive Officer

Road(s) and routes per MDRS Map: Cow Dung road, Galileo road (and Watney road)

Mode of travel: Walking & driving

GreenHab Report – April 4th

Report title: GreenHab Report
Crew #: 333
Position: GreenHab Officer
Report prepared by: Matthias De Groote
Date: 04-04-2026
Sol: 6

Environmental control (fan & heater): Yes, Fan max: 80, Fan min: 75; Heater: 70 F°

Average temperatures (last 24h): 83.25 F°
Maximum temperature (last 24h): 97.9 F°
Minimum temperature (last 24h): 68.6 F°

Hours of supplemental light: 4:00

Daily water usage for crops: 2.25 gal
Daily water usage for research and/or other purposes: 0.5 gal
Water in the Tank (160 gal useful capacity): 144 gallons remaining
Time(s) of watering for crops: 10:00 am

Changes to crops: Crops were attacked, basil, radishes and lettuce were severely impacted. Loss has been estimated to 1/3 of radishes, 70% of the basil bush (more than 60g of leaves) and the lettuce head (80-90% of the lettuce).

A net over the crops and a cage mouse trap have been installed to try and protect the crops.

If any attack of this sort has been reported in the past, please inform us of how it’s been dealt with.

Harvest: N/A

Support/supplies needed: No

Supplemental Operations Report – April 4th

Date: 04/04/2026
Name of person filing report: Brett Bennett
Reason for Report: Routine
Off-Nominal Systems: Hab outer shell. GreenHab hydrant. Tunnel tarps.

Power system: Nominal.
Current generator hours: 1448.0

Propane:
Station Tank: 52%
Director Tank: 68%
Intern Tank: 70%
Generator Tank: 76%

Water:
Hab Static Tank: 360 gallons
GreenHab: 145 gallons
Outpost tank: 450 gallons
Septic Tanks emptied: No
The water transfer pump in the Hab car showed reduced flow (approx 25% normal) on Apr 1. Inspection showed a functioning pump and damaged pump hose. A new hose has been ordered and the pump fitted with a set of interim hoses.
Continue to refer to last week’s operations report for the most recent information on the GreenHab hydrant.

Rovers:
Sojourner rover used: No (except for servicing)
Hours: 226.7
Beginning Charge: 100%
Ending Charge: 100%
Currently Charging: Yes
Monthly rover service for all rovers completed April 4. All rovers are in a nominal condition.

Cars:
Hab Car used and why, where: Multiple trips to Hanksville for supplies.
Crew Car used and why, where: Replacement vehicle is still being sourced.
General notes and comments: The station will be operating with the hab car only until crew car function is replaced.

Internet: Nominal.
EVA suits and radios: Nominal. Recent issue with headset wire connection fixed by crew.
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus wide inspection, if action taken, what and why: Several tunnel tarps show wind damage from an April 2 storm. Tarps have been temporarily repaired by crew 333 and are functional but will ultimately need replacement.
The RAM personnel door was also damaged during the April 2 storm, showing a 0.5" overlap on the latch side post-storm. A disassembly and repair of the top hinge has restored the door to nominal function.
Hab outer shell shows ongoing wear.

Summary of Hab Operations: Nominal. See notes on recent additional projects in preceding reports.

Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Nominal.
Summary of Outpost Operations:
Director trailer: Nominal. Trailer holding tank emptied and water refilled on April 4.
Intern trailer: Nominal. Water refilled on April 4.

Journalist Report – April 3rd

Report title: Journalist Report
Crew #: 333
Position: Crew Journalist
Report prepared by: Joanna Galloway
Date: 03-04-2026
Sol: 5

Journalist Report Title: Sol 5

The weather finally calmed down today therefore everyone was able to go out in EVA. Joanna, Antoine and Matias were up first, after a hefty bowl of oatmeal they made their way to sea of shells to continue Antoine’s experience. Being worried about the batteries of the rovers, they didn’t go all the way to Sea of Shells but they still got to appreciate some beautiful views along the way. Marie, Matthias and Zahraa used their morning to work on some homework and to set up their experiments.

Lunch was minestrone soup ‘à la martienne’ but before we could attack it, we had to do an emergency EVA to repair the pressurised tunnels that led to the RAM. The wind must’ve made it collapse. Matias and Antoine were there for the rescue, repairing the issues, this took quite some time as the job was much bigger than we initially thought. They first thought they could seal the hole but quickly found out that the tarp will have to be completely replaced. Lunch was eating quickly to get the second team ready for their afternoon EVA.

Marie, Matthias and Zahraa completed Antoine’s experience before heading to Sea of Shells. Quite a surreal site to see on Mars. They climbed multiple hills to take in the beautiful scenery from the highest point. The rest of the team back at the station, cleaned the Hab and passed the time by playing games. They also enjoyed an interview of the Artemis Crew who are currently in space.

When the Crew was reunited, we made the most of the time by writing the reports and planning the next few days. For dinner we are having chicken curry prepared by Zahraa and Matthias. Tonight will be an evening playing mime games and an early night to prepare for tomorrow’s big day.

Astronomy Report – April 3rd

[category  astronomy-report]

Report title: Astronomy Report
Crew #: 333
Position: Crew Astronomer
Report prepared by: Matias Ballivian Ascarrunz
Date: 03-04-2026
Sol:5

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: MDRS-WF and RCOS-16
Objects to be Imaged this Evening: NGC-7000, Messier 86, Messier 51
Images submitted with this report: None
Problems Encountered: None

MUSK OBSERVATORY
Solar Features Observed: N/A
Images submitted with this report: N/A
Problems Encountered: N/A